Calculate the energy of an electron in the n = 7 level of a hydrogen atom.


1
Expert's answer
2021-02-09T03:44:09-0500

E = "\frac{13.6}{n^2}" × eV


E = energy , n = 7 , eV = 1.6×10-19


E = "\frac{13.6}{7^2}" × 1.6×10-19 J


E = "\frac{13.6}{49}" × 1.6×10-19 J


E = 0.444×10-19 J